OpenStudy (anonymous):

how do I find the average rate of change of a function, y=x^3 + 1 over interval [2,3]?

myininaya (myininaya):

\[AR=\frac{f(t_1)-f(t_2)}{t_1-t_2}\]

myininaya (myininaya):

x=t

myininaya (myininaya):

\[AR=\frac{(3^3+1)-(2^3+1)}{3-2}\]

OpenStudy (anonymous):

THanks very much!

myininaya (myininaya):

np

OpenStudy (anonymous):

so if the function is g(x)=x^2 between [-1,1], the AR= 0? Do I have that right?

OpenStudy (anonymous):

is average rate of change of the function, g(x)=x^2, over interval [-1,1] equal to 0?

myininaya (myininaya):

yes